### v3

# The Discrepancy of Shortest Paths

Abstract
The hereditary discrepancy of a set system is a certain quantitative measure of the pseudorandom properties of the system. Roughly, hereditary discrepancy measures how well one can $2$-color the elements of the system so that each set contains approximately the same number of elements of each color. Hereditary discrepancy has well-studied applications e.g. in communication complexity and derandomization. More recently, the hereditary discrepancy of set systems of shortest paths has found applications in differential privacy [Chen et al. SODA 23].
The contribution of this paper is to improve the upper and lower bounds on the hereditary discrepancy of set systems of unique shortest paths in graphs. In particular, we show that any system of unique shortest paths in an undirected weighted graph has hereditary discrepancy $\widetilde{O}(n^{1/4})$, and we construct lower bound examples demonstrating that this bound is tight up to hidden polylog $\text{polylog }n$ factors. Our lower bounds apply even in the planar and bipartite settings, and they improve on a previous lower bound of $\Omega(n^{1/6})$ obtained by applying the trace bound of Chazelle and Lvov [SoCG’00] to a classical point-line system of Erdős. We also show similar bounds on (non-hereditary) discrepancy and in the setting of directed graphs.
As applications, we improve the lower bound on the additive error for differentially-private all pairs shortest distances from $\Omega(n^{1/6})$ [Chen et al. SODA 23] to $\widetilde{\Omega}(n^{1/4})$, and we improve the lower bound on additive error for the differentially-private all sets range queries problem to $\widetilde{\Omega}(n^{1/4})$, which is tight up to hidden polylog $\text{polylog }n$ factors [Deng et al. WADS 23].
Appendix A Related Work
A.1 Discrepancy on graphs
In graph theory, the discrepancy of a graph introduced by Erdős [erdHos1963combinatorial] is defined as follows:
$\max_{U\subset V}\left|e(G[U])-p{{|U|}\choose 2}\right|,$
where $e(G[U])$ is the number of edges of the induced subgraph $G[U]$ on vertices $U$ and $p=|E|/{n\choose 2}$ is the density of edges. If we consider a complete graph and randomly color each edge with probability $p$, the above definition of discrepancy quantifies the deviation of induced subgraphs of $G$ from their expected size. Erdős and Spencer [erdiis1972imbalances] showed that the graph discrepancy is $\Theta(n^{3/2})$ when $p=1/2$. This definition and related definitions (e.g., positive discrepancy, dropping the absolute operator) have applications to quasi-randomness [chung1989quasi], graph cuts and edge expansion [alon1997edge, bollobas2006discrepancy, raty2023positive]. There is also study of multicolor discrepancy [gishboliner2022discrepancies, freschi2021note] that we skip here.
Of particular relevance to our work, Balogh et al. [balogh2020discrepancies] studied edge discrepancy (as defined in this paper) of (spanning) trees, paths and Hamilton cycles of a graph $G$. In particular, they showed that, for any labeling of edges of $G$, there is a path with discrepancy $\Omega(n)$, even when the graph is a grid. Prior to this, either probabilistic construction exhibiting such a lower bound was known [beck1983size, dudek2015alternative, letzter2016path] or an explicit construction of linear size non-planar graphs [alon1988explicit], resolving a question by Erdős [erdHos1963combinatorial]. The construction for the planar graph in [balogh2020discrepancies] can be extended to coloring of vertices such that there is a path with vertex discrepancy $\Omega(n)$ when there are exponentially many paths and $\Omega(\sqrt{n})$ when there are polynomially many paths. (see Appendix D for details).
Discrepany of paths in directed graphs has also been studied. Reimer [reimer2002ramsey] showed that, if a directed graph has discrepancy $\Omega(n)$, then the graph must have $\Omega(n^{2})$ edges. In the case when we do not allow antiparallel edges, Ben-Eliezer et al. [ben2012size] showed that there is a directed graph with $\Theta(n^{2}\log^{2}(n))$ edges such that any mapping $\chi:E\to\{-1,1\}$ will either have a path of length $\Omega(n)$ and all edges mapped to $-1$ or a path of length $\Omega(n\log(n))$ with all edges mapped to $+1$.
A.2 Connection with curve discrepancy
A classical topic in computational geometry is to study upper and lower bounds of the discrepancy/hereditary discrepancy of the incidence matrix of geometric objects and a set of points. For example, for a set of $n$ points and $n$ halfplanes in $\mathbb{R}^{2}$, the $n$ by $n$ incidence matrix (with rows corresponding to halfplanes and columns corresponding to points) has discrepancy of $\Omega(n^{1/4})$. For $n$ points and $n$ lines in the plane, the discrepancy of the incidence matrix is $\Omega(n^{1/6})$ [Chazelle2000-ld]. In general the discrepancy of such incidence matrix is related to the ‘complexity’ of the geometric shapes. In our setting of a graph, the set of all pairs shortest paths defines a set system on the vertices. When the graph is planar, the shortest paths are essentially simple curves in the plane.
We would like to compare our results with discrepancy of curves and points in the geometric setting. Using the classification in [Pach1998-ib], a family of simple curves have $k$ degree of freedom and multiplicity type $s$ if for any $k$ points there are at most $s$ curves passing through all of them, and any pair of curves intersect in at most $s$ points. Lines in the plane have degree of $2$ and multiplicity of $1$. A set of curves with degree of $2$ and multiplicity of $1$ is called pseudolines – two pseudolines have at most one intersection. For $n$ points and $n$ lines, the discrepancy is upper bounded by $O(n^{1/6}(\log n)^{2/3})$ [Chazelle2000-rd] which is nearly tight by a polylogarithmic factor. The proof uses the standard partial coloring argument with the Szeremédi-Trotter bound on point-line incidence – for any $n$ points and $m$ lines there are at most $O(m^{2/3}n^{2/3}+m+n)$ point-line incidences [Szemeredi1983-tv]. The Szeremédi-Trotter bound can be extended to a set of pseudolines [Pach1998-ib, Szekely1997-yp]. Therefore the same proof and upper bound hold for the discrepancy of pseudolines.
For a consistent set of shortest paths, two shortest paths will only intersect at a contiguous segment, which may have multiple vertices/points. Thus using the curve classification criterion, a consistent family of shortest paths in the plane has degree of $2$ but multiplicity $s$ that is possibly higher than a constant. In fact, our discrepancy lower bound construction in the planar graph setting uses a design with $s$ possibly as high as $n^{1/2}$. This is the major difference of shortest paths in a planar graph with pseudolines, which allows the discrepancy of shortest paths to go beyond the pseudoline upper bound of $\widetilde{O}(n^{1/6})$ [^1] Using the incidence upper bound for $k=2$ and $s=n^{1/2}$ from [Szekely1997-yp] and partial coloring, one can obtain a discrepancy upper bound of $\widetilde{O}(n^{1/3})$ for our path construction. In contrast, we obtained nearly tight bound of $\widetilde{\Theta}(n^{1/4})$..
Appendix B Technical Preliminaries
Technical Observations for Discrepancy on Consistent Path Systems
We first observe that, using a random coloring $\chi$, we can guarantee that for all paths $\pi\in\Pi$ [Chazelle2000-rd]:
$|\chi(\pi)|\leq\sqrt{2|\pi|\ln(4|\Pi|)}.$
This immediately provides a few observations.
Observation B.1.
When $\Pi$ is a set of paths with size polynomial in $n$, then $\operatorname{disc}(\Pi)=O(\sqrt{n\log n})$. This bound is true even for paths that are possibly non-consistent.
Observation B.2.
When the longest path in $\Pi$ has $D$ vertices we have $\operatorname{disc}(\Pi)=O(\sqrt{D\log n})$. Thus, for graphs that have a small diameter (e.g., small world graphs), the discrepancy of shortest paths is automatically small.
We have defined vertex and edge (hereditary) discrepancy; one may wonder if there is an underlying relationship between vertex and edge (hereditary) discrepancy since they share the same bounds in most scenarios as presented in LABEL:tab:res-summary. The following observation shows that vertex discrepancy directly implies bounds on edge discrepancy.
Observation B.3.
Denote by $\operatorname{disc}(n)$ (and $\operatorname{herdisc}(n)$) the maximum discrepancy (minimum hereditary discrepancy, respectively) of a consistent path system of a (undirected or directed) graph of $n$ vertices. We have that
  1. 1.
    Let $g(x)$ be a non-decreasing function. If $\textnormal{herdisc}_{v}(n)\geq g(n)$, then $\textnormal{herdisc}_{e}(n)\geq g(n/2)$.
  2. 2.
    Let $f(x)$ be a non-decreasing function. If $\textnormal{disc}_{v}(n)\leq f(n)$, then $\textnormal{disc}_{e}(m)\leq f(m)$.
Proof.
We prove Item 1 by showing that if graph $G=(V,E,w)$ with the consistent path matrix $A^{(v)}$ has hereditary discrepancy at least $g(n)$, we can obtain another graph $G^{\prime}=(V^{\prime},E^{\prime},w^{\prime})$ and matrix $A^{(e)}$ as the (consistent path) edge incidence matrix with hereditary discrepancy at least $g(n/2)$. The construction is as follows.
  1. (a)
    We first split each vertex $v\in V$ in $G$ to two vertices $(v_{in},v_{out})$ to obtain $V^{\prime}$.
  2. (b)
    For every $v\in V$, add a single edge $(v_{in},v_{out})$ to $E^{\prime}$.
  3. (c)
    For any $v\in V$ and each edge $(u,v)\in E$ (with the fixed $v$), add edges $(u_{out},v_{in})$ and $(u_{in},v_{out})$ to $E^{\prime}$.
The path incident matrix $A^{(e)}$ is defined as follows: for each path as a row $a$ of $A^{(v)}$, construct a new path in $G^{\prime}$ by following the order of $u_{out}\rightarrow v_{in}\rightarrow v_{out}\rightarrow w_{in}$ for a $u\rightarrow v\rightarrow w$ sequence. For each row in $A^{(v)}$, we mark the used edges as $1$ in $A^{(e)}$ with the path constructed by the above process. Note that the new path system defined by $A^{(e)}$ remains consistent: for any intersection between the two paths $P_{1}\cap P_{2}=(u_{1},u_{2},\cdots,u_{\ell})$, the intersection remains a single path of $(u_{1,in},u_{1,out},\cdots,u_{\ell,in},u_{\ell,out})$ in $A^{(e)}$.
Let $Y$ be the columns that induces the $g(n)$ discrepancy on $G$, i.e.,
$\min_{x\in\{-1,+1\}^{\left|Y\right|}}\|A^{(v)}_{Y}x\|_{\infty}=g(n).$
Now, observe that, for each row in $A^{(e)}$, an edge $(v_{in},v_{out})$ is marked as $1$ if and only if $v$ is marked as $1$ in $A^{(v)}$. Therefore, we can take the new set $Y^{\prime}$ as the edges corresponding to $Y$, and there is
$\min_{x\in\{-1,+1\}^{\left|Y^{\prime}\right|}}\|A^{(e)}_{Y^{\prime}}x\|_{\infty}=\min_{x\in\{-1,+1\}^{\left|Y\right|}}\|A^{(v)}_{Y}x\|_{\infty}=g(n).$
Finally, since graph $G^{\prime}$ has $n^{\prime}=2n$ vertices, we have the hereditary discrepancy to be at least $g(n^{\prime}/2)$, as desired.
For Item 2, we show that the hereditary edge discrepancy of $G$ is at most $f(m)$, which implies the discrepancy upper bound. For a graph with $n$ vertices, $m$ edges, and a path incident matrix $A^{(e)}$, suppose $Y$ is the set of columns (edges) that attain the hereditary discrepancy. We can add a vertex $v_{e}$ for each $e\in Y$ and construct a new path incident matrix $A^{(v)}$, which is a matrix with $\left|Y\right|$ rows. Concretely, for each row of $A^{(v)}$, we simply let vertices $v_{e}\in Y$ be $1$ if the corresponding edge is used in $A^{(e)}$. By the consistency of $A^{(e)}$, the new path incident matrix also characterizes a consistent path system (we can think of the underlying graph as the complete graph on vertex set of $Y$). Note that since there are at most $m$ vertices in $A^{(e)}$, we can get $f(m)$ discrepancy for the path system characterized by $A^{(e)}$. This implies a $f(m)$ hereditary edge discrepancy on the original path system, which in turn implies the desired discrepancy upper bound.
Finally, note that the argument remains valid when the graph is directed, which means the results hold for both undirected and directed graphs. ∎
Known Results in Discrepancy Theory
The first result that we discuss is the one that gives an upper bound on discrepancy of a set system in terms of primal shatter function.
Definition 1 (Primal Shatter Function).
Let $(X,\mathcal{R})$ be a set system, i.e., $X$ is a ground state and $S=\{S_{1},S_{2},\cdots,S_{\ell}\}$ with $S_{i}\subseteq X$ for all $1\leq i\leq\ell$. Let $s$ be a positive integer. The primal shatter function, denoted as $\pi_{\mathcal{R}}(s)$, is defined as $\max_{A\subseteq X:\,|A|=s}\left|\{A\cap S\mid S\in\mathcal{R}\}\right|$.
The following is a well known result in discrepancy theory.
Proposition B.4 (Theorem 1.2 in [Matousek1995-as]).
Given a set system $(X,\mathcal{R})$, the discrepancy of a range space $\mathcal{R}$ whose primal shatter function is bounded by $\pi_{\mathcal{R}}(x)=cx^{d}$, for some constant $c>0$, $d>1$, is
$O(n^{1/2-1/(2d)}),$
where $n$ is the size of the ground state, and $O(\cdot)$ hides the dependency on $c$ and $d$.
For the lower bound on hereditrary discrepancy, one general result is the trace bound [Chazelle2000-ld].
Lemma B.5 (Trace Bound [Chazelle2000-ld]).
If $A$ is an $m$ by $n$ incidence matrix and $M=A^{T}A$, then
$\operatorname{herdisc}(A)\geq\frac{1}{4}c^{n\cdot\operatorname{tr}(M^{2})/(\operatorname{tr}(M))^{2}}\sqrt{\frac{\operatorname{tr}(M)}{n}},$
where $10<c<$ 1 is a constant.
Recently, this trace bound has been improved in a series of works culminating in the following bound in Larsen [larsen2017constructive] which we also use in our work:
Lemma B.6 (Trace Bound [larsen2017constructive]).
If $A$ is an $m$ by $n$ incidence matrix and $M=A^{T}A$, then
$\operatorname{herdisc}(A)\geq{(\operatorname{tr}(M))^{2}\over 8e\operatorname{tr}(M^{2})\cdot\min\{m,n\}}\sqrt{\frac{\operatorname{tr}(M)}{\max\{m,n\}}}.$
We give various interpretations of $\operatorname{tr}(M)$ and $\operatorname{tr}(M^{2})$ in Lemma B.6 that would be useful later on. Algebraically, $\operatorname{tr}(M)$ is the sum of its eigenvalue while $\operatorname{tr}(M^{2})$ is the sum of square of the eigenvalues. Combinatorially, $\operatorname{tr}(M)$ is the number of ones in $A$ and $\operatorname{tr}(M^{2})$ is the number of rectangles of all ones in $A$. Geometrically, $\operatorname{tr}(M)$ is the count of point/region incidences and $\operatorname{tr}(M^{2})$ is number of pairs of points in all the pairwise intersections of regions. Finally, if $A$ is the incidence matrix for shortest path, $\operatorname{tr}(M^{2})$ is the number of length $4$ cycles in the underlying graph. Based on the algebraic interpretation, it means that the trace bound is non-trivial whenever all the eigenvalues of $A$ are fairly uniform. This can be seen by noticing that, if $\{\lambda_{1},\cdots,\lambda_{n}\}$ are eigenvalues of $A$, then
${\operatorname{tr}(M)^{2}\over\operatorname{tr}(M)}=n\cos^{2}(\theta),$
where $\theta$ is the angle between the vector $(\lambda_{1},\cdots,\lambda_{n})$ and the all one-vector.
Concentration Inequalities
We use the following standard variants of the Chernoff-Hoeffding bound in our paper.
Proposition B.7 (Chernoff bound).
Let $X_{1},\ldots,X_{n}$ be $n$ independent random variables with support on $\{0,1\}$. Define $X:=\sum_{i=1}^{n}X_{i}$. Then, for every $\delta>0$, there is
$\Pr\left(X\geq(1+\delta)\cdot\mathbb{E}\left[X\right]\right)\leq\exp\left(-\frac{\delta^{2}}{\delta+2}\cdot\mathbb{E}\left[X\right]\right).$
In particular, when $\delta\in(0,1]$, there is
$\Pr\left(\left|X-\mathbb{E}\left[X\right]\right|>\delta\cdot\mathbb{E}\left[X\right]\right)\leq 2\cdot\exp\left(-\frac{\delta^{2}\mathbb{E}\left[X\right]}{3}\right).$
Proposition B.8 (Additive Chernoff bound).
Let $X_{1},\ldots,X_{n}$ be $n$ independent random variables with support in $[0,1]$. Define $X:=\sum_{i=1}^{n}X_{i}$. Then, for every $t>0$,
$\Pr\left(\left|X-\mathbb{E}\left[X\right]\right|>t\right)\leq 2\cdot\exp\left(-\frac{2t^{2}}{n}\right).$
Appendix C Missing Proofs from LABEL:subsec:edge-disc-exp
Proof of LABEL:prop:disc_E_1.
The proposition follow from an argument similar to Proposition LABEL:prop:disc_bound. For each path $\pi^{\prime}\in\Pi^{\prime}$, let $X_{\pi^{\prime}}$ be the random variable defined as
$X_{\pi^{\prime}}:=\left|\sum_{e\in E(\pi)\cap E(\pi^{\prime})}\chi(e)\right|.$
We can restate the discrepancy of $E_{1}$ as
$\left|\sum_{e\in E_{1}}\chi(e)\right|=\sum_{\pi^{\prime}\in\Pi^{\prime}}X_{\pi^{\prime}}.$
By Proposition LABEL:prop:edge_small_disc, if $|E(\pi)\cap E(\pi^{\prime})|\equiv 0\mod 2$, then $X_{\pi^{\prime}}=0$, so without any loss of generality, we may assume that $|E(\pi)\cap E(\pi^{\prime})|$ is odd for all $\pi^{\prime}\in\Pi^{\prime}$. In this case,
$\Pr\left[X_{\pi^{\prime}}=-1\right]=\Pr\left[X_{\pi^{\prime}}=1\right]=1/2,$
implying that $\mathbb{E}[\sum_{\pi^{\prime}\in\Pi^{\prime}}X_{\pi^{\prime}}]=0$. Then by Proposition LABEL:prop:edge_path_cover and the Chernoff bound, it follows that for any constant $c\geq 1$,
$\Pr\left[\sum_{\pi^{\prime}\in\Pi^{\prime}}X_{\pi^{\prime}}\geq c\cdot n^{1/4}\log^{1/2}n\right]\leq e^{-c^{2}\frac{n^{1/2}\log n}{2|\Pi^{\prime}|}}\leq e^{-c^{2}/2\cdot\log n}\leq n^{-c^{2}/2}.$
Proof of LABEL:prop:disc_E_2.
The proposition follow from an argument similar to Proposition LABEL:prop:disc_bound. Note that by the Edge Covering Property of the path cover (Proposition LABEL:prop:edge_path_cover),
$|E_{2}|=\left|\{(u,v)\in E(\pi)\mid u,v\not\in V^{\prime}\}\right|\leq n^{1/2}.$
Moreover, the edges in $E\setminus E^{\prime}$ are labeled independently at random, so we may apply a Chernoff bound to argue that for any constant $c\geq 1$,
$\Pr\left[\left|\sum_{e\in E_{2}}\chi(e)\right|\geq c\cdot n^{1/4}\log^{1/2}n\right]\leq e^{-c^{2}\frac{n^{1/2}\log n}{2|E_{2}|}}\leq e^{-c^{2}/2\cdot\log n}\leq n^{-c^{2}/2}.$
completing the proof. ∎
Proof of LABEL:prop:disc_E_3.
Let
$k:=\left|\{\pi^{\prime}\in\Pi^{\prime}\mid\pi\cap\pi^{\prime}\neq\emptyset\}\right|$
denote the number of paths in our path cover that intersect $\pi$. We define a function $f:\mathbb{Z}_{\geq 0}\mapsto\mathbb{Z}_{\geq 0}$ such that $f(\phi)$ equals the largest possible value of $|E_{3}|$ when $\phi=k$. Note that $f$ is well-defined since $0\leq|E_{3}|\leq|E|$. We will prove that $f(\phi)\leq 4\phi,$ by recursively decomposing path $\pi$. When $\phi=1$, there is only one path $\pi^{\prime}\in\Pi^{\prime}$ that intersects $\pi$. Then the only edges in $E_{3}$ are of the form
$E(\pi)\cap((V^{\prime}\times(V\setminus V^{\prime}))\cup((V\setminus V^{\prime})\times V^{\prime}))=E(\pi)\cap((\pi^{\prime}\times(V\setminus\pi^{\prime}))\cup((V\setminus\pi^{\prime})\times\pi^{\prime})).$
By the Consistency Property of Proposition LABEL:prop:edge_path_cover, path $\pi^{\prime}$ can intersect $\pi$ and then split apart at most once. Then
$f(1)=|E_{3}|=|E(\pi)\cap((\pi^{\prime}\times(V\setminus\pi^{\prime}))\cup((V\setminus\pi^{\prime})\times\pi^{\prime}))|\leq 2.$
When $\phi>1$, we will split our analysis into the two cases:
  • Case 1. There exists paths $\pi_{1}^{\prime},\pi_{2}^{\prime}\in\Pi^{\prime}$ and nodes $v_{1},v_{2},v_{3}\in\pi$ such that $v_{1},v_{3}\in\pi_{1}^{\prime}$ and $v_{2}\in\pi_{2}^{\prime}$ and $v_{1}<_{\pi}v_{2}<_{\pi}v_{3}$. In this case, we can assume without any loss of generality that $\pi[v_{1},v_{3}]\cap\pi_{1}^{\prime}=\{v_{1},v_{3}\}$ (e.g., by choosing $v_{1},v_{3}$ so that this equality holds). Let $x$ be the node immediately following $v_{1}$ in $\pi$, and let $y$ be the node immediately preceding $v_{3}$ in $\pi$. Recall that $s$ is the first node of $\pi$ and $t$ is the last node of $\pi$. It will be useful for the analysis to split $\pi$ into three subpaths:
    $\pi=\pi[s,v_{1}]\circ\pi[x,y]\circ\pi[v_{3},t],$
    where $\circ$ denotes the concatenation operation. Let
    $\phi_{1}:=|\{\pi^{\prime}\in\Pi^{\prime}\mid\pi[x,y]\cap\pi^{\prime}\neq\emptyset\}|\text{\qquad and \qquad}\phi_{2}:=|\{\pi^{\prime}\in\Pi^{\prime}\mid(\pi[s,v_{1}]\circ\pi[v_{3},t])\cap\pi^{\prime}\neq\emptyset\}|.$
    We claim that $\phi_{1}<\phi$, $\phi_{2}<\phi$, and $\phi_{1}+\phi_{2}=\phi$. We will use these facts to establish a recurrence relation for $f$. By our assumption that $\pi[v_{1},v_{3}]\cap\pi_{1}^{\prime}=\{v_{1},v_{3}\}$, it follows that $\pi[x,y]\cap\pi_{1}^{\prime}=\emptyset$, and so $\phi_{1}<\phi$. Likewise, by the No Repeats Property of Proposition LABEL:prop:edge_path_cover,
    $(\pi[s,v_{1}]\circ\pi[v_{3},t])\cap\pi_{2}^{\prime}=\emptyset,$
    so $\phi_{2}<\phi$. Finally, observe that more generally, if there exists a path $\pi^{\prime}\in\Pi^{\prime}$ such that $\pi^{\prime}\cap\pi[x,y]\neq\emptyset$ and $\pi^{\prime}\cap(\pi[s,v_{1}]\circ\pi[v_{3},t])\neq\emptyset$, then the No Repeats Property of Proposition LABEL:prop:edge_path_cover is violated. We conclude that $\phi_{1}+\phi_{2}=\phi$.
    Now $|E_{3}|$ can be upper bounded by the following inequality:
    $|E_{3}|\leq|E_{3}\cap E(\pi[x,y])|+|E_{3}\cap E(\pi[s,v_{1}]\circ\pi[v_{3},t])|+2.$
    Then using the observations about $\phi_{1},\phi_{2}$, and $\phi$ in the previous paragraph, we obtain the following recurrence for $f$:
    $f(\phi)\leq f(\phi_{1})+f(\phi_{2})+2=f(i)+f(\phi-i)+2,$
    where $0<i<\phi$.
  • Case 2. There exists a path $\pi^{\prime}\in\Pi^{\prime}$ and $v_{1},v_{2}\in\pi$ such that $\pi\cap\pi^{\prime}=\pi[v_{1},v_{2}]\cap V^{\prime}$. Let $x$ be the node immediately preceding $v_{1}$ in $\pi$, and let $y$ be the node immediately following $v_{2}$ in $\pi$. Again, we split $\pi$ into three subpaths:
    $\pi[s,t]=\pi[s,x]\circ\pi[v_{1},v_{2}]\circ\pi[y,t].$
    Let
    $\phi_{1}:=|\{\pi^{\prime}\in\Pi^{\prime}\mid\pi[v_{1},v_{2}]\cap\pi^{\prime}\neq\emptyset\}|\text{\qquad and \qquad}\phi_{2}:=|\{\pi^{\prime}\in\Pi^{\prime}\mid(\pi[s,x]\circ\pi[y,t])\cap\pi^{\prime}\neq\emptyset\}|.$
    By our assumption in Case 2, it follows that $\phi_{1}=1$ and $\phi_{2}=\phi-1$. Since $|E_{3}|$ can be upper bounded by the inequality
    $|E_{3}|\leq|E_{3}\cap E(\pi[v_{1},v_{2}])|+|E_{3}\cap E(\pi[s,x]\circ\pi[y,t])|+2,$
    we immediately obtain the recurrence
    $f(\phi)\leq f(\phi_{1})+f(\phi_{2})+2\leq f(1)+f(\phi-1)+2.$
Taking our results from Case 1 and Case 2 together, we obtain the recurrence relation
$f(\phi)\leq\begin{cases}\max\left\{f(i)+f(\phi-i)+2,\text{ }f(1)+f(\phi-1)+2\right\}&\phi>1\text{ and }1<i<\phi\\ 2&\phi=1\end{cases}$
Applying this recurrence $\leq\phi$ times, we find that
$f(\phi)\leq\phi\cdot f(1)+2\phi\leq 4\phi.$
Finally, since $k\leq|\Pi^{\prime}|\leq n^{1/2}$ and we defined $f$ so that $f(k)$ equals the largest possible value of $|E_{3}|$, we conclude that
$|E_{3}|\leq f(k)\leq f(n^{1/2})=O(n^{1/2}).$
Since the edges in $E_{3}\subseteq E\setminus E^{\prime}$ are labeled independently at random, we may apply a Chernoff bound as in Proposition LABEL:prop:edge_disc_e2 to argue that $\chi(E_{3})=O(n^{1/4}\log^{1/2}n)$ with high probability. ∎
Appendix D Discrepancy Bounds for Paths Without Consistency
In a graph when we consider simple paths without the consistency requirement, there is a strong lower bound on both vertex and edge discrepancy. In particular, we have the following theorem.
Theorem 1.
There is a planar graph $G=(V,E)$ such that the following is true for any coloring $f:V\mapsto\{-1,1\}$ of vertices $V$:
  1. 1.
    There is a family $\Pi$ of simple paths with $|\Pi|=O(\exp(n))$ and vertex discrepancy of $\Omega(n)$.
  2. 2.
    There is a family $\Pi$ of simple paths with $|\Pi|=O(n)$ and vertex discrepancy of $\Omega(\sqrt{n})$.
The same claim holds true for edge discrepancy as well.
Proof.
The first claim follows from Proposition 1.6 in [balogh2020discrepancies], which says that the edge discrepancy of paths on a $k\times\ell$ grid graph is at least $\Omega(k\ell)$. To prove vertex discrepancy, we make two additional remarks about this construction. First, for our purpose, it is sufficient to consider only an $n\times 2$ grid graph $G$. The set of paths in the construction of [balogh2020discrepancies] consists of all paths that start from the top left corner and bottom left corner going to the right and possibly taking a subset of the vertical edges in the grid graph. The number of paths is $O(2^{n})$. Second, for vertex discrepancy, we define a companion graph $G^{\prime}$. Specifically, for each grid edge $e$ in $G$, we place a vertex $v$ of $G^{\prime}$ on $e$. We connect two vertices in $G^{\prime}$ if and only if the corresponding edges in $G$ share a common vertex. The graph $G^{\prime}$ is still planar. In additional, a path $P$ in $G$ maps to a corresponding path $P^{\prime}$ in $G^{\prime}$ where vertices on $P^{\prime}$ follow the same order of the corresponding edges on $P$. See Figure 1 for an example. Therefore the edge discrepancy in $G$ and the vertex discrepancy of $G^{\prime}$ are the same.
Figure 1: A $n\times 2$ grid graph (with vertices shown in hollow and edges in black) $G$ with one path (in blue) starting from the top left corner go the right. The solid vertices and edges in dashed red define the companion graph $G^{\prime}$. The corresponding path in $G^{\prime}$ is shown in pink.
For the second claim, we take an $n\times n$ Hadamard matrix $H$ with $n$ as power of $2$. The elements in $H$ are $+1$ or $-1$. Each row has $n/2$ elements of $+1$ and $n/2$ elements of $-1$ and the rows are pairwise orthogonal. It is known [Chazelle2000-rd] that the matrix $A=\{a_{ij}\}=\frac{1}{2}(H+J)$ with $J$ as an $n\times n$ matrix of all $1$ has discrepancy at least $\Omega(\sqrt{n})$.
Now we try to embed the matrix $A$ by paths on a $n\times 2$ grid graph $G$. Denote by $e_{j}$ as the $j$th vertical edge in $G$, $1\leq j\leq n$. For the $i$th row of $A$, we define set $X_{i}=\{e_{j}:a_{ij}=1\}$. We then define two paths $P(X_{i})$, $P^{\prime}(X_{i})$ on $G$.
  • Path $P(X_{i})$ starts from the top left corner of $G$ going to the right and the vertical edges $P(X_{i})$ visits are precisely $X_{i}$.
  • Path $P^{\prime}(X_{i})$ starts from the bottom left corner of $G$ going to the right and, similar to $P(X_{i})$, the vertical edges $P(X_{i})$ visits are precisely $X_{i}$.
Note that $P(X_{i})$ and $P^{\prime}(X_{i})$ each contains edges $X_{i}$ (see Figure 2 for an illustration). Also $P(X_{i})$ and $P^{\prime}(X_{i})$ do not share any horizontal edges, and collectively cover all horizontal edges in $G$. In addition, we define two paths $P$ and $P^{\prime}$ with $P$ starting from the top left corner and visiting all the top horizontal edges and $P^{\prime}$ starting from the bottom left corner and visiting all bottom horizontal edges. We have $22n+$ 2 paths in total – each row $i$ of the Hadamard matrix contributes $2$ paths, and we additionally use $P$ and $P^{\prime}$.
Figure 2: A $n\times 2$ grid graph $G$ with two paths $P(X_{i})$ and $P^{\prime}(X_{i})$. The $i$th row of matrix $A$, i.e., $A_{i}=(\frac{1}{2}(H+J))_{i}$, corresponds to $X_{i}=(1,1,0,1,1,1,0,0)$.
For any {±$\{\pm 1\}$ coloring $f$ on edges in the grid graph $G$, define $D$ as the maximum absolute value of the sum of the colors of edges of each path, among all the $22n+$ 2 paths. We now argue that $D=\Omega(\sqrt{n})$.
First we define an $n$-dimensional vector $\chi\in\{+1,-1\}^{n}$ with $\chi_{j}=f(e_{j})$, i.e., the color of the $j$th vertical edge in $G$. Since the discrepancy of matrix $A$ is $\Omega(\sqrt{n})$, there must be one row vector $a_{i}$ in $A$ such that $a_{i}\cdot\chi\geq c\sqrt{n}$ for some constant $c$. In other words, the sum of the colors of the edges in $X_{i}$ is $x=\sum_{e_{j}\in X_{i}}f(e_{j})$ with $|x|\geq c\sqrt{n}$. Without loss of generality, we assume $x>0$, which in turn implies $x\geq c\sqrt{n}$.
Now we define
$z_{1}:=\sum_{e\in P(X_{i})\setminus X_{i}}f(e)\quad\text{ and}\quad z_{2}:=\sum_{e\in P^{\prime}(X_{i})\setminus X_{i}}f(e).$
If $\max\{z_{1},z_{2}\}\geq 0$, then we are done as the total sum of colors of either $P(X)$ or $P^{\prime}(X)$ is at least $x\geq c\sqrt{n}$. Otherwise, suppose we have $z_{1}<0$ and $z_{2}<0$. We now have from path $P(X_{i})$, $D\geq x+z_{1}$, and from path $P^{\prime}(X_{i})$, $D\geq x+z_{2}$. Further, consider paths $P$ and $P^{\prime}$, the total sum of colors of all the horizontal edges is $z_{1}+z_{2}$. Since both $z_{1}$ and $z_{2}$ are negative, there should be
$2D\geq\sum_{e\in P}f(e)+\sum_{e\in P^{\prime}}f(e)=-z_{1}-z_{2}.$
Summing up all three inequalities, we have $4D\geq 2x$. Thus $D\geq x/2\geq c\sqrt{n}/2$. This finishes the proof for edge discrepancy, and the vertex discrepancy bound can be obtained using the same trick as in the proof of claim 1. ∎
The above theorem provides lower bounds for discrepancy. Since hereditary discrepancy is at least as high as discrepancy, the lower bounds hold for hereditary discrepancy as well.
We also want to remark that the paths used in the above theorem are $2$-approximate shortest paths. The shortest path from top left corner to the top right corner of the $n\times 2$ grid is of length $n-1$ while all paths used are of length at most $22n-$ 2. The shortest path from top left corner to bottom right corner is of length $n$ and all paths used in the construction are of length at most $12n-$ 1. Therefore when we relax from shortest paths to $2$-approximate shortest paths the discrepancy bounds substantially go up. If we allow weights on the edges we can make these paths to be $1+\varepsilon$-approximate shortest paths for any $\varepsilon>0$.
Grid graphs are a special family of planar graphs. Actually for grid graphs we can say a bit more on discrepancy of shortest paths. If we take shortest paths on an unweighted grid graph (without even requiring consistency property and there could be exponentially many shortest path between two vertices), the discrepancy is $O(1)$. Specifically, for vertex discrepancy on a $k\times\ell$ grid graph of left bottom corner at the origin and the top right corner at coordinate $(k-1,\ell-1)$, if we give a color of $+1$ to all vertices of coordinate $(x,y)$ with even $x+y$ and a color of $-1$ to all other vertices, any shortest path visits a sequence of vertices with sum of coordinates alternating between even and odd values and thus has a total color of $O(1)$. For edge discrepancy, for all horizontal edges we give color $+1$ ($-1$) if the left endpoint is at an even (odd) $x$-coordinate and the right endpoint is at an odd (even) $x$-coordinate. We do the same for vertical edges. Again any shortest path has a ‘staircase’ shape and a total coloring of $O(1)$.
Appendix E Relation Between Discrepancy and Hereditary Discrepancy in General Graphs
In Theorem LABEL:thm:undi-hdisc-lb, we showed a construction of a path weighted graph $G$ whose system of unique shortest paths $\Pi$ satisfies $\operatorname{herdisc}_{v}(\Pi)\geq\Omega(n^{1/4}/\sqrt{\log(n)})$. Here, we will observe that this result extends to discrepancy:
Theorem 2.
There are examples of $n$-node undirected weighted graphs $G$ with a unique shortest path between each pair of nodes in which this system of shortest paths $\Pi$ has
$\textnormal{disc}_{v}(\Pi)\geq\Omega\left({n^{1/4}\over\sqrt{\log(n)}}\right).$
Proof.
Let $G$ be the graph from Theorem LABEL:thm:undi-hdisc-lb, and let $\Pi$ be its system of unique shortest paths. By definition of hereditary discrepancy, there exists an induced path subsystem $\Pi^{\prime}\subseteq\Pi$ with
$\operatorname{disc}_{v}(\Pi^{\prime})\geq\Omega\left({n^{1/4}\over\sqrt{\log(n)}}\right).$
Recall that, by induced subsystem, we mean that we may view the paths of $\Pi$ as abstract sequences of nodes, and then $\Pi^{\prime}$ is obtained from $\Pi$ by deleting zero or more nodes and deleting all occurrences of those nodes from the middle of paths. Thus the paths in $\Pi^{\prime}$ are not still paths in $\Pi$, but they are paths in a different graph $G^{\prime}$ on $n^{\prime}\leq n$ nodes. It thus suffices to argue that all paths in $\Pi^{\prime}$ are
It thus suffices to argue that there is a graph $G^{\prime}$ on $n^{\prime}\leq n$ nodes in which all paths in $\Pi^{\prime}$ are unique shortest paths. Indeed, this is well known, and is shown e.g. in [Bodwin19] (c.f. Lemma 2.4.4 and 2.4.11). To sketch the proof: suppose that a node $v$ is deleted from the initial system $\Pi$. Consider each path $\pi\in\Pi$ that contains $v$ as an internal node, i.e., it has the form
$\pi=(\dots,u,v,x,\dots).$
When $v$ is removed, the path now contains the nodes $u,x$ consecutively, and so we must add $(u,x)$ as a new edge to $G^{\prime}$ so that $\pi$ is a path in $G^{\prime}$. We judiciously set the edge weight to be $w(u,x):=w(u,v)+w(v,x)$. The weighted length of the path $\pi$ does not change, and yet the distances of $G^{\prime}$ majorize those of $G$, which implies that $\pi$ is still a unique shortest path in $G^{\prime}$. Inducting this analysis over each deleted vertex leads to the desired claim. ∎
Appendix F Application in Matrix Analysis
Hereditary discrepancy is intrinsically related to factorization norm, which has found applications in many ares of computer science, including but not limited to quantum channel capacity, communication complexity, etc. For any complex matrix $A\in\mathbb{C}^{m\times n}$, its factorization norm, denoted by $\gamma_{2}(A)$, is defined as the following optimization problem:
$\gamma_{2}(A)=\min\left\{\|L\|_{2\to\infty}\|R\|_{1\to 2}:A=LR\right\}.$ (1)
Here, for $p,q\in\mathbb{N}_{\geq 0}$,
$\|{A}\|_{p\to q}=\max\frac{\|Ax\|_{q}}{\|x\|_{p}}.$
One can write eq 1 in a form of a semi-definite program (see Lee et al. [lee2008direct]) and also show that the Slater point exists. In particular, the primal and dual program coincides. An interesting question in matrix analysis is to estimate the factorization norm of different class of matrices. In a series of work, various authors have computed tight bounds on the factorization norm of certain class of matrices:
  • If $A$ is a unitary matrix, then $\gamma_{2}(A)=1$ [schur1911remarks].
  • If $A\in\mathbb{R}^{n\times n}$ is a positive semidefinite matrix with entries $A_{ij}$ [schur1911remarks], then
    $\gamma_{2}(A)=\max_{1\leq i\leq n}A_{ii}.$
  • Mathias [mathias1993hadamard] $A$ satisfies the property that $\sqrt{A^{\top}A}\bullet\mathbb{I}=\sqrt{AA^{\top}}\bullet\mathbb{I}={\operatorname{tr}(A)\over n}\mathbb{I}$, then
    $\gamma_{2}(A)={\mathsf{Tr}(A^{\top}A)\over n}.$
    In particular, if $A\in\{0,1\}^{n}$ is a lower-triangular one $1$ matrix, then $\gamma_{2}(A)=\Theta(\log n)$ [fichtenberger2023constant, henzinger2023almost, kwapien1970main].
  • If $A\in\mathbb{R}^{n\times n}$ is a lower-triangular Toeplitz matrix with entries decreasing either polynomially or exponentially, then $\gamma_{2}(A)=\Theta(1)$ [henzinger2023unifying].
Our tight bound on hereditary discrepancy for consistent path on graphs allows us to give tight bound on the factorization norm for the corresponding incidence matrix. In particular, we use the following result:
Lemma F.1 (Matoušek et al. [matouvsek2020factorization]).
For any real $m\times n$ matrix $A\in\mathbb{R}^{m\times n}$, there exists absolute constants $0<c<C$ such that
$c\frac{\gamma_{2}(A)}{\sqrt{\log(m)}}\leq\operatorname{herdisc}(A)\leq C\gamma_{2}(A)\log(m).$
Combining Lemma F.1 with our results, we have the following corollary:
Corollary F.2.
Let $A_{G}$ be the incident matrix for unique shortest path system on an $n$ vertices graph $G$. Then if $G$ is bipartite, planar, or any general graph, then $\gamma_{2}(A_{G})=\widetilde{\Theta}(n^{1/4})$.